# A Light-Cone Approach to Higher-Order Cosmological Observables

### v1

@@ Acknowledgements
We are very thankful to Ruth Durrer, Matteo Magi and Fabian Schmidt for useful discussions. PB and GM are supported in part by the Istituto Nazionale di Fisica Nucleare (INFN) through the Commissione Scientifica Nazionale 4 (CSN4), under the Iniziativa Specifica (IS) Theoretical Astroparticle Physics (TAsP) and the IS Quantum Fields in Gravity, Cosmology and Black Holes (FLAG). The work of GM and MRMS was supported by the research grant number 2022E2J4RK ”PANTHEON: Perspectives in Astroparticle and Neutrino THEory with Old and New messengers” under the program PRIN 2022 funded by the Italian Ministero dell’Università e della Ricerca (MUR) and by the European Union – Next Generation EU. GF acknowledges the COST Action CosmoVerse, CA21136, supported by COST (European Cooperation in Science and Technology). GF is also member of the Gruppo Nazionale per la Fisica Matematica (GNFM) of the Istituto Nazionale di Alta Matematica (INdAM).

### v2

@@ Acknowledgements
We are very thankful to Ruth Durrer, Matteo Magi and Fabian Schmidt for useful discussions. PB and GM are supported in part by the Istituto Nazionale di Fisica Nucleare (INFN) through the Commissione Scientifica Nazionale 4 (CSN4), under the Iniziativa Specifica (IS) Theoretical Astroparticle Physics (TAsP) and the IS Quantum Fields in Gravity, Cosmology and Black Holes (FLAG). The work of GM and MRMS was supported by the research grant number 2022E2J4RK “PANTHEON: Perspectives in Astroparticle and Neutrino THEory with Old and New messengers” under the program PRIN 2022 funded by the Italian Ministero dell’Università e della Ricerca (MUR) and by the European Union – Next Generation EU. GF acknowledges the COST Action CosmoVerse, CA21136, supported by COST (European Cooperation in Science and Technology). GF is also member of the Gruppo Nazionale per la Fisica Matematica (GNFM) of the Istituto Nazionale di Alta Matematica (INdAM). The work of MRMS was supported by the Brazilian National Council for Scientific and Technological Development – CNPq, under project grant 447129/2024-4.
@@ Appendix A Free-Falling Observer
Appendix B Table of Notation
[…]
Table 1: List of the main symbols used in this paper. Here, $\mathbb{S}^{2}$ is the 2-dimensional sphere, PG stands for “Poisson Gauge”, $P_{\text{o}}$ is an observer monopole term and $v_{||\text{o}}$ is the radial peculiar velocity at the observer position.